# High stakes exams inflate gender gap thus leading to systematic grading errors in introductory physics

### v1

@@ IV.3 Highest stakes assessments
Since 203 of the classes in our database have final exam scores it seems useful to examine those data because the final exam scores contribute much more to the course grade than any in-term exam and so can be considered to be high-stakes exams. Unfortunately we don’t have a low-stakes version of the final exam to compare with the high-stakes versions. However, we do have the medium-stakes in-term exam scores in each class which should be good predictors of final exam scores in that class because they are written and graded by the same instructors and largely cover the same material. In fact, we find a moderate/strong correlation between the in-term exam average and the final exam score (correlation coefficient = 0.65, $p<10^{-3}$) but a smaller correlation between student GPA and either in-term or final exam score (both correlation coefficients are about 0.56, $p<10^{-3}$). We’ve already suggested that the medium-stakes in-term exam scores include a bias against female students, so it may be that using these scores as predictors of final exam scores will not only control for the demonstrable knowledge/ability of each student in the particular course but may also control for any gender bias inherent in exams. If this were the case then the controlled final exam gender gap should be approximately zero. To test this idea we will use HLM to fit the following model
@@
The resulting coefficients from the HLM fit to Eq. 3 are given in Table 3. The most notable coefficient is $b_{Female}$ and the most notable thing about this coefficient is that it is quite clearly not zero. In fact, the (in-term exam controlled) final exam gender gap is half the size of the gender gap for the in-term exams alone. This tells us that the full gender gap on the high-stakes final exam is about 50% larger than that of the medium-stakes in-term exams. So the high-stakes final exam has a much larger gender gap than the medium-stakes in-term exams which, themselves, have a much larger gender gap than the low-stakes first-attempt in-term exams in the retake classes. And $b_{Female*Retake}$ is consistent with there being a similar (in-term-exam controlled) gender gap in the high-stakes final exam in the retake classes. Again, it seems likely that the importance (or weight) of the exam (the exam stakes) is closely related to the size of the grade penalty received by female students.
Table 3: The coefficients from an HLM fit to final exam z-scores (equation 3) are shown along with their standard errors, z-statistics, and p-values. Included are 203 classes that had a total of 49,085 students. The interaction term suggests that the gender gap is not significantly different for the retake classes.
@@ V Discussion
In our introduction we present many student-level variables that may contribute to the gender gap in introductory physics. But these factors; preparation, interest, anxiety, stereotypy threat, self-efficacy, identity, and belonging are all deeply intertwined with both larger influences of society, and the cultural context of the physics discipline. For instance, results from Leslie et. al. [24] on the gender imbalance in academic fields (including physics) that are felt to require “innate talent” rather than persistent effort has led these researchers as well as others (Muradoglu et. al. [25]) to argue that, in fields perceived to require “innate talent” women (and other groups that are often intellectually marginalized) have more negative experiences, feel less welcome, have more “imposter” feelings, and are more vulnerable to stereotype threat [26]. This general idea can connect the various gender gap explanations that we have listed above. We postulate that many of these feelings of otherness are exaggerated in a high-pressure environment. With this stated assumption we can then argue that the reason the gender gap exists is not because of all these negative biases against women, it’s because these biases exist AND we put women in a high pressure environment.
@@
Furthermore, an extremely important takeaway from this logic is that tests are not measuring what we think they are measuring because if they were, performance gaps between demographic groups would not depend on the stakes of the particular exams as Sec. IV.3 shows they do. The implications are that if we care about measuring physics knowledge and skills, we need to be looking at ways to reduce or eliminate the stakes associated with our assessments because since the stakes of a given timed in-person assessment correlate with the size of the gender gap, exams are currently not good measures of physics knowledge and skills across demographic groups.
@@
Another problem with a focus on the individual student level issues that large groups of students may be having in physics is that this leads too naturally to a consideration of those groups of students as deficient or inadequate in some way. Researchers tend to use a Student Deficit model rather than the idea that these students are just reacting, on average, the way anyone would if they were part of a group stereotyped by society. Indeed, results from Salehi et al [8] fit this picture - they find that the gender gaps for exams are much larger in general science and engineering courses where women are less represented compared to biological science courses where women are more represented and may therefore feel more like they belong in that space.
@@
Cotner and Ballen [2] cut the Gordian knot made up of interconnected possible student deficits by suggesting the easiest thing to do is to simply change the structure of the course. They proposed using a Course Deficit model rather than a Student Deficit model. Our recent work [1, 27] supports the idea that changing the course structure is enough to eliminate equity gaps. While studies that examine the effects that society (and more specifically disciplinary cultures) have on different demographic groups are important and necessary, we should not present those effects as the reasons for the differences between the performances of demographic groups. If we move in that direction we miss the forest for the trees. For example, the solution to the gender gap in physics exams isn’t to make women feel less anxious in a stressful environment, it’s to change the environment to be less stressful. This is the power that we have as instructors. We might not be able to change the impact society has on women in physics, but we can mitigate the effects of those impacts in our courses.
Instructors reading this paper might consider adoption of retake exams or other assessments that lower the stakes associated with testing in introductory physics courses.
VI Conclusions
@@
Building from our prior research on this dataset, [1] where we found that retake exams were associated with a zero gender gap in introductory courses, we rule out the possibility that this happens because women are more likely to take advantage of the retake opportunity, as men are just as likely to take advantage of the retakes. We also find that the reason for eliminated gender gap doesn’t appear to be that women need an extra chance to fail and catch up (presumably due to any differences in incoming physics experiences) because women appear to do as well on the first try of a retake exam as the men.
@@
Instead we find that women do as well as and perhaps slightly better (though not significantly so) than men on the first try of a retake than their male peers. We also find that women do significantly better overall on all of the exams (independently true for both first and second tries) in the retake course than they do in the courses without retake exams. Finally, if the stakes of the final exam didn’t matter for the gender gap, we would find that controlling for in-term exam scores (an assessment of the same type, but not the same stakes) would zero the gender gap on the final exams. Instead we find that controlling for in-term exam scores on the final exams in courses without retakes, does not remove the gender gap in final exams, again suggesting that the stakes of the exam are responsible for the gender gap in physics.
@@ Appendix B Including a student only once doesn’t change our numbers much
As we point out in Sec. III, many of the students in our dataset are represented more than once because they were enrolled in, and received grades in, more than one class included in the data. One worry we had was that including these students more than once might influence our numbers or our conclusions. To check this we have, for each of the 26,766 unique students in the datafile, randomly chosen only one of their classes to include in an analysis. We made 10 of these random choices in order to check the coefficients and errors in Table 2. We find that the model coefficients we present in the text do not change significantly and that the error estimates change as $\sqrt{1/N}$ , where $N$ is the total number of sets of student grades, as one expects if one simply had fewer data points. So the results and conclusions we report in the text are essentially unchanged by having a set of students present in the data more than once because they took more than one class in the series.
Table 4: The average coefficients from 10 HLM fits to equation 2 (where each fit corresponds to a different random choice of which one class, for each student, is included) are shown along with their standard errors, z-statistics, and p-values. The comparison with Table 2, which includes all classes of all students, shows that the only important effect is to increase the error estimates by the square root of ratio of number of students included in the fit. For this table the average number of students included (because we know their GPA’s) averages 24,641. And the retake classes include only 368 students instead of the 57,806 students and 396 students, respectively, included in Table 2.
@@ Appendix C Results from the intersections of gender and race/ethnicity
In Fig. 3 we show the results for $b_{Female*Retake}$, as determined from a fit to Eq. 2 for each fit to the particular racial/ethnic group described in Table 5. One sees that these differences are all positive although most of them are not significantly different than zero. However, it is notable that several are well above zero.

### v2

@@ IV.3 Highest stakes assessments
Since 203 of the classes in our database have final exam scores it seems useful to examine those data because the final exam scores contribute much more to the course grade than any in-term exam and so can be considered to be high-stakes exams. Unfortunately we don’t have a low-stakes version of the final exam to compare with the high-stakes versions because even in the courses with retake options on the exams, it was not possible to retake the final exam. However, we do have the medium-stakes in-term exam scores in each class which should be good predictors of final exam scores in that class because they are written and graded by the same instructors and largely cover the same material. In fact, we find a moderate/strong correlation between the in-term exam average and the final exam score (correlation coefficient = 0.65, $p<10^{-3}$) but a smaller correlation between student GPA and either in-term or final exam score (both correlation coefficients are about 0.56, $p<10^{-3}$). We’ve already suggested that the medium-stakes in-term exam scores include a bias against female students, so it may be that using these scores as predictors of final exam scores will not only control for the demonstrable knowledge/ability of each student in the particular course but may also control for any gender bias inherent in exams. If this were the case then the controlled final exam gender gap should be approximately zero. To test this idea we will use HLM to fit the following model
@@
The resulting coefficients from the HLM fit to Eq. 3 are given in Table 3. The most notable coefficient is $b_{Female}$ and the most notable thing about this coefficient is that it is quite clearly not zero. In fact, the (in-term exam controlled) final exam gender gap is half the size of the gender gap for the in-term exams alone. This tells us that the full gender gap on the high-stakes final exam is about 50% larger than that of the medium-stakes in-term exams. So the high-stakes final exam has a much larger gender gap than the medium-stakes in-term exams which, themselves, have a much larger gender gap than the low-stakes first-attempt in-term exams in the retake classes. And $b_{Female*Retake}$ is consistent with there being a similar (in-term-exam controlled) gender gap in the high-stakes final exam in the retake classes. Again, it seems likely that the importance (or weight) of the exam (the exam stakes) is closely related to the size of the grade penalty received by female students. Finally, as noted in Appendix B, we have checked that these coefficients and the resulting conclusions hold when we (randomly) include each student only once in our analyses.
Table 3: The coefficients from an HLM fit to final exam z-scores (equation 3) are shown along with their standard errors, z-statistics, and p-values. Included are 203 classes that had a total of 48,428 students. The interaction term suggests that the gender gap is not significantly different for the retake classes.
@@ V Discussion
In our introduction we present many student-level variables that may contribute to the gender gap in introductory physics. But these factors; preparation, interest, anxiety, stereotypy threat, self-efficacy, identity, and belonging are all deeply intertwined with both larger influences of society, and the cultural context of the physics discipline. For instance, results from Leslie et al. [22] on the gender imbalance in academic fields (including physics) that are felt to require “innate talent” rather than persistent effort has led these researchers as well as others (Muradoglu et. al. [23]) to argue that, in fields perceived to require “innate talent” women (and other groups that are often intellectually marginalized) have more negative experiences, feel less welcome, have more “imposter” feelings, and are more vulnerable to stereotype threat [32]. This general idea can connect the various gender gap explanations that we have listed above. We postulate that many of these feelings of otherness are exaggerated in a high-pressure environment. With this stated assumption we can then argue that the reason the exam gender gap exists is not because of all these negative biases against women, it’s because these biases exist AND we put students in a high pressure environment.
@@
Furthermore, an extremely important takeaway from this logic is that tests are not measuring what we think they are measuring because if they were, performance gaps between demographic groups would not depend on the stakes of the particular exams as Sections IV.2 and IV.3 show they do. These two sections provide two independent results, each supporting the idea that increasing the stakes of an exam increases the gender gap in a way that is disconnected from differences in physics knowledge/skills. The implications are that if we care about measuring physics knowledge and skills, we need to be looking at ways to reduce or eliminate the stakes associated with our assessments. Since the stakes of a given timed in-person assessment correlate with the size of the gender gap, exams are currently not good measures of physics knowledge and skills across demographic groups.
@@
Another problem with putting one’s focus on the individual student level issues that large groups of students may be having in physics is that this leads too naturally to a consideration of those groups of students as deficient or inadequate in some way. Researchers tend to use a Student Deficit model rather than the idea that these students are just reacting, on average, the way anyone would if they were part of a group stereotyped by society. Indeed, results from Salehi et al [13] fit this picture - they find that the gender gaps for exams are much larger in general science and engineering courses where women are less represented compared to biological science courses where women are more represented and may therefore feel more like they belong in that space.
@@
Cotner and Ballen [2] cut the Gordian knot made up of interconnected possible student deficits by suggesting the easiest thing to do is to simply change the structure of the course. They proposed using a Course Deficit model rather than a Student Deficit model. Our recent work [1, 33] supports the idea that changing the course structure is enough to eliminate equity gaps. While studies that examine the effects that society (and, more specifically, disciplinary cultures) have on different demographic groups are important and necessary, we should not present those effects as the reasons for the differences between the performances of demographic groups. If we move in that direction we miss the forest for the trees. Perhaps the most important take-away from our paper is that the solution to decreasing the gender gap in physics exams isn’t to make women feel less anxious in a stressful environment, it’s to change the environment to be less stressful. This is the power that we have as instructors. We might not be able to change the impact society has on women in physics, but we can mitigate the effects of those impacts in our courses. The implications of this work indicate that instructors should strongly consider adopting assessments such as retakes that lower the stakes associated with testing in introductory physics courses. Timed in-person exams are typically anxiety inducing experiences, however our work shows that if we change the stakes of these exams by allowing a retake, that women perform as well as men. This suggests that the stakes are influencing the outcome of exams so that the exams are inaccurate measures of differences between men and women. Instructors who want measurements of physics understanding less influenced by stakes (or anxiety) should adopt assessments with lower stakes.
Our discovery that gender gaps are dependent on the level of stakes associated with timed exams has an extremely important implication for education researchers. It is common for researchers to use prior exam results (e.g. quantitative SAT, ACT, & Advanced Placement exams) as a proxy for preparation for physics. Women’s scores on these exams tend to be lower than men’s [34] (This was some of the impetus for proposing research question #2.) Therefore, a common resulting finding from those studies is that women’s past preparation for physics is somehow less than men’s and that this is what is driving the gender gap. An alternative interpretation of the same numerical results is that women are experiencing the same issue in both contexts, and that high-stakes exams are causing the gender gap, instead of measuring it. Therefore, this study adds evidence to our previous work indicating that it may not make sense to use exam scores as a proxy for preparation when comparing across group demographics [1]. We suggest that future research using exam scores as a proxy for preparation proceed with thoughtful caution.
VI Limitations
All of the retake exam classes (as well as the non-retake classes) were offered within the CLASP (Collaborative Learning through Active Sense-making in Physics) series of courses [35, 33]. These courses are studio type courses where students spend almost 80% of their class time working in discussion/lab sections, generally in groups of five. The effects of allowing retake exams may well be different in classes that are predominantly lecture-based. In addition, almost the entire course grade in the courses studied was determined by the exam grades in these courses. It could be that in a course where most of the grade is determined by timed in-person solo assessments, differences of the impact of the stakes of each of those assessments are easier to identify. We encourage replication studies of this phenomenon.
VII Conclusions
@@
Building from our prior research on this dataset [1], where we found that retake exams were associated with a zero gender gap in introductory courses, we rule out the possibility that this happens because women are more likely to take advantage of the retake opportunity, as men are just as likely to take advantage of the retakes. We also find that the reason for the elimination of the gender gap doesn’t appear to be that women need an extra chance to fail and catch up via practice or extra help (presumably due to any differences in incoming physics preparation) because women appear to do as well as men on their (low-stakes) first exam attempt in the retake classes. These negative results are already important findings because they provide counter evidence to a recurring theme in the literature that gender gaps are due to differences in preparation. The women in the retake classes did not need extra time nor extra tries to perform equally to the men in those classes and so showed no obvious deficiencies.
@@
Instead we find that women do as well as and perhaps slightly better (though not significantly so) than men on the first, low stakes, exam in a retake class. We also find that women do noticeably better overall on each of the exams (independently true after both first and second tries) in the retake course than they do in the courses without retake exams.
Our data on final exams provide a second, and independent, measurement of an exam gender gap that is inflated by the stakes of the exam. If the stakes of the final exam didn’t matter for the gender gap, we would find that controlling for in-term exam scores (assessments of the same type, but not the same stakes) would zero the gender gap on the final exams. Instead we find that controlling for in-term exam scores on the final exams in courses with or without retakes, does not remove the gender gap in final exams, again suggesting that the stakes of the exam control the size of the gender gap in introductory physics exams. The size of our data set allows us to measure the extra gender gap on the final exam in these classes to better than one part in ten.
Instructors interested in removing gender performance differences related to exam stakes should reduce the stakes of their assessments by using lower stakes assessment strategies such as implementing retakes on exams.
VIII Acknowledgements
The authors would like to thank the National Science Foundation for their partial support of this project through the now terminated NSF HSI IUSE 1953760 grant. This work would not have been possible without the support of San José State University’s Research, Scholarship, and Creative Activity Assigned Time Program.
@@ Appendix B Including a student only once doesn’t change our numbers much
As we point out in Sec. III, many of the students in our dataset are represented more than once because they were enrolled in, and received grades in, more than one class included in the data. One worry we had was that including these students more than once might influence our numbers or our conclusions. To check this we have, for each of the 26,783 unique students in the datafile, randomly chosen only one of their classes to include in an analysis. We made 10 of these random choices in order to check the coefficients and errors in Table 2. The coefficient resulting from fits with those ten choices were then averaged and are shown in Table 4. We find that the model coefficients we present in the text do not change significantly and that the error estimates change as $\sqrt{1/N}$ , where $N$ is the total number of sets of student grades, as one expects if one simply had fewer data points. So the results and conclusions we report in the text are essentially unchanged by having a set of students present in the data more than once because they took more than one class in the series.
Our conclusions from the numerical analysis of final exam gender gaps given in Section IV.3 are also unchanged by including each student only once. When we randomly choose only one appearance for each student in this final exam analysis we find that the gender gap is that same as shown in Table 3 but that the error estimate on the gender gap increases to 0.011 from 0.007 (so z-statistic to -10.9 from -16.2).
Table 4: The average coefficients from 10 HLM fits to equation 2 (where each fit corresponds to a different random choice of which one class, for each student, is included) are shown along with their standard errors, z-statistics, and p-values. The comparison with Table 2, which includes all classes of all students, shows that the only important effect is to increase the error estimates by the square root of ratio of number of students included in the fit. For this table the average number of students included (because we know their GPA’s) is 24,641 and the retake classes include only 368 students instead of the 57,806 students and 396 students, respectively, included in Table 2.
@@ Appendix C Results from the intersections of gender and race/ethnicity
In Fig. 3 we show the results for $b_{Female*Retake}$, as determined from a fit to Eq. 2 for each particular racial/ethnic group described in Table 5. One sees that these differences are all positive although most of them are not significantly different than zero. However, it is notable that several are well above zero.
@@ Appendix D Grade scale effect on gender gap
Appendix E Retake exams gender gap
The group of students taking a retake exam are a self-chosen subset of the full set of students and that subset changes from exam to exam with the exam averages depending on who decides to retake an exam. For these reasons, it is probably unwise to draw any conclusions from analyses of these data. Nevertheless, for completeness we will calculate the gender gap on the retake exams which we have considered to be medium stakes exams. Since the group of students retaking an exam changes from week to week we will just find the gender gap for each week’s retake exam and then take a weighted average with the weights determined by the square of the standard error of that exam’s gender gap. We, again, z-score each retake exam in each class and use HLM to determine the gender gap after controlling for GPA as in Eq. 1. We find that the weighted average of those four retake exam gender gaps is -0.06 $\pm$ 0.08. This number is certainly consistent with zero gap but the 95% confidence interval also overlaps with the other medium stakes gaps shown in Table 2. It probably isn’t surprising that we have no conclusions to draw from these calculations.
Appendix F Students who are Non-binary or for whom we don’t have gender data
Since the university didn’t collect gender self-identity data until after essentially all of the exams in our dataset had been taken there is very little we can say regarding students who don’t identify as one of the two genders that most populate our data. Nevertheless, there were 5 self-identified non-binary students in our data as well as well as 31 students whose genders unknown by the university. None of the nonbinary students appeared in a retake class. For completeness we will extend the final exam analysis we did in Section IV.3 to include both non-binary students and students whose gender was unknown to the university. We define a categorical variable Gender = 0 for students listed as male in our dataset, =1 for those listed as female, =2 for those listed as non-binary, and =3 for those whose gender is unknown. We extend our final exam analysis to include Gender by using HLM to fit
$\begin{split}FinalExam&=b_{0}+b_{InTrmExm}InTrmExm\\ &+b_{Gender}Gender\end{split}$ (4)
where the coefficients of $Gender$ each represent the average in-term-exam-controlled final exam difference between that group and the group of male students. All of the coefficients from the HLM fit to Eq. 4 are given in Table 6. The large p-values for the Non-Binary and Unknown genders suggest that, as expected, there aren’t enough data to measure the differences from the male students on final exam scores.
Table 6: The coefficients from an HLM fit to final exam z-scores (equation 4) are shown along with their standard errors, z-statistics, and p-values. Included are 203 classes that had a total of 48,478 students. The coefficients corresponding to the different genders each represent the average difference from the group of male students.
Coeff. Value Error z-statistic p-value
$b_{InTrmExm}$ 0.9652 0.0051 187.4 $<10^{-3}$
$b_{Female}$0.1157 0.0071 -16.3 $<10^{-3}$
$b_{NonBinary}$ 0.27 0.28 0.96 034
$b_{Unknown}$0.06 0.12 -0.50 0.62
$b_{0}$ 0.0744 0.0056 13.3 $<10^{-3}$